OpenStudy (anonymous):

what is the vertex of y=0.5x^2-2x-4 ?

OpenStudy (anonymous):

LETS MAKE BABIES!!!

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

I SAID LETS MAKE SOME LOVEEEEE

OpenStudy (anonymous):

noooo im trying to do homework

OpenStudy (anonymous):

AFTER U R DONE WITH TAT

OpenStudy (anonymous):

-4,0 IS THE ANSWER BTW

OpenStudy (aum):

y=0.5x^2-2x-4 ---- (1) The coefficients are: a = 0.5, b = -2, c = -4 The x-coordinate of the vertex is given by: -b / 2a = -(-2) / ( 2 * 0.5) = 2 / 1 = 2 To find the y-coordinate of the vertex, put x = 2 in equation (1) and find y: y = 0.5 * (2)^2 - 2(2) - 4 = 2 - 4 - 4 = -6 Vertex is at (2, -6).